# Evaluating Scoring Bias in LLM-as-a-Judge

### v3

@@ Abstract
The remarkable capabilities of Large Language Models (LLMs) have led to the emergence of “LLM-as-a-Judge”, a paradigm where LLMs serve as automated evaluators for complex tasks. This approach has been widely adopted across diverse fields, including natural language processing and preference learning. However, the fairness and reliability of these LLM judges are often compromised by inherent biases. While current research has predominantly focused on biases in comparison-based evaluations, systematic investigations into scoring-based evaluations remain limited. To address this gap, we first formally define scoring bias as the discrepancy in scores when the judge model’s prompt is exposed to bias-inducing perturbations. We then propose a comprehensive framework to evaluate this bias. This framework is supported by two key contributions: (1) new evaluation datasets constructed by augmenting existing benchmarks with synthesized data, and (2) a suite of multi-faceted metrics designed for scoring bias analysis. Our experiments reveal that the scoring stability of LLM judges is significantly disrupted by these biases. Furthermore, our findings offer actionable insights for designing more robust scoring prompts and mitigating biases related to score rubrics, score IDs, and reference answer selection.
Introduction
Large Language Models (LLMs), with outstanding performance, have been widely applied to Natural Language Generation (NLG) tasks such as summarization (Zhang et al. 2024), question answering (Robinson, Rytting, and Wingate 2022), and domain-specific dialogues (Wang et al. 2024a; Singhal et al. 2025). In these tasks, the model’s responses require fine-grained evaluation of various aspects such as quality, factuality, and instruction-following, thereby providing valuable insights for model improvement. Traditional statistical NLG evaluation metrics, such as BLEU (Papineni et al. 2002) and ROUGE (Lin 2004), evaluate the generated text by measuring n-gram overlap between the text and the golden answer. The above methods exhibit two primary limitations: (1) They focus solely on the similarity between the generated text and the golden answer, rather than quantifying intrinsic attributes of the text such as writing style, factuality, and educational value. (2) The resulting metrics lack interpretability, making it difficult to provide guidance for the model improvement. There is an urgent need for fine-grained, multi-faceted, and interpretable evaluation metrics and methods tailored to NLG tasks and LLM applications. With remarkable capabilities in instruction understanding and response generation, LLMs can provide diverse forms of evaluation output, such as scoring, comparison, and ranking. In addition, they can conduct evaluations along specified dimensions and generate critiques for their judgments, thus enhancing the interpretability of the evaluation. This line of research, referred to as “LLM-as-a-Judge” (Gu et al. 2024; Li et al. 2024a), has recently emerged as a prominent focus.
@@ Introduction
Robustness is a critical factor in ensuring the fairness and reliability of LLM judgments, which requires that, under identical judge rubrics, LLM judges should provide consistent judgments for semantically equivalent but differently phrased evaluation targets. However, recent work (Park et al. 2024) indicates that LLM judges are subject to various biases, which undermine the self-consistency of their judgments. Many studies have sought to identify the types of bias that cause shifts in judgments. Wang et al. finds that the judgments of the candidate responses can be easily influenced by altering their position in the context. Wataoka, Takahashi, and Ri measures the self-preference bias in LLM-as-a-Judge, which refers to the tendency of LLM judges to overestimate the quality of their own outputs. Wei et al. defines length bias as the tendency of LLM judges to favor longer or shorter responses.
Although existing studies have systematically investigated various biases in LLM-as-a-Judge, several pressing issues remain to be addressed. Firstly, while some work (Wang et al. 2024b; Ye et al. 2024a) has addressed biases such as positional bias and refinement awareness in scoring-based LLM judges, many existing studies remain focused on biases in comparative LLM judges. Existing work (Ye et al. 2024a) attempts to investigate specific biases in scoring tasks, such as self-enhancement bias and refinement-aware bias. However, there is still a lack of comprehensive research on biases in scoring-based LLM-as-a-Judge. Secondly, the biases defined in existing work are predominantly related to the choice of response models and the attributes or presentation forms of the evaluation targets, such as position and length. However, variation inherent in the LLM judge itself, such as the ordering of score rubrics and the selection of reference answers, may also impact the self-consistency of judgments, which lacks a systematic investigation in current research.
To solve the above problems, we propose scoring bias, defined as the scoring shifts of LLM-as-a-Judge under perturbations of the scoring prompt, as Figure 1 shows. In this work, we propose three types of scoring bias: score rubric order bias, score ID bias, and reference answer score bias, and design a framework to systematically evaluate these biases. We select four existing LLM-as-a-Judge benchmarks as evaluation datasets and utilize correlation with golden scores and score distributions as metrics to quantify the robustness and scoring tendency of LLM judges. Our key contributions can be summarized as follows:
  • We conduct the first systematic evaluation of scoring bias in LLM-as-a-Judge, introducing a formal definition of this bias and a comprehensive methodology to quantify it from multiple perspectives.
  • To address the critical lack of datasets for scoring bias analysis, we introduce an automatic data synthesis pipeline that expands existing LLM-as-a-Judge benchmarks into a tailored evaluation corpus, thereby establishing a foundational resource for future studies.
  • We empirically demonstrate that even the most advanced LLMs suffer from substantial scoring biases and exhibit obvious self-bias tendency. Furthermore, our analysis yields valuable recommendations for designing more robust scoring prompts and mitigating scoring biases.
Related Work
LLM-as-a-Judge
Gu et al. defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context. Common forms of LLM judgment include score generation (Gao et al. 2023), pairwise comparison (Qin et al. 2024), ranking (Li, Patel, and Du 2023), multiple-choice selection (Li et al. 2024b), and critique generation (Ke et al. 2024). Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts (Liu et al. 2023), improvement of LLM’s abilities such as fine-tuning (Wang et al. 2024c), multi-LLM collaboration pipelines such as feedback mechanisms (Xu et al. 2023), and output optimization strategies such as multi-output ensembling (Sottana et al. 2023). The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement (Thakur et al. 2024), Cohen’s Kappa, and Spearman’s correlation (Bai et al. 2023). LLM-as-a-Judge is widely applied to evaluation (Lin and Chen 2023), alignment (Bai et al. 2022), retrieval (Sun et al. 2023), reasoning (Gao et al. 2024a), and other specific domains (Ryu et al. 2023; Babaei and Giudici 2024). LLM-as-a-Judge faces challenges of bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges (Li et al. 2024a).
Bias in LLM-as-a-Judge
Ye et al. defines the prompt for LLM judges as three components: system instruction, question, and response. Based on this, they define bias in LLM-as-a-Judge as the tendency for judgments to differ when the system instruction or responses are modified in a bias-related manner. Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc. These biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks. (Zheng et al. 2023; Wei et al. 2024) Currently, the types of bias remain a focal point of investigation, and numerous benchmarks have been proposed to evaluate bias in LLM-as-a-Judge. OffsetBias (Park et al. 2024) identifies six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on. CALM (Ye et al. 2024a) quantifies and analyzes 12 types of bias in LLM judges. However, the biases defined in current work are predominantly focused on comparative evaluation tasks, and there is a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks. Similar to research on adversarial attacks, the evaluation metrics for bias in LLM-as-a-Judge are often based on the consistency between bias-related modified judgments and the original judgments, or on alignment with golden answers.
Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges. Liu et al. introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators. Wang et al. proposes three methods to calibrate positional bias in LLM-as-a-Judge, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration. Our evaluation results on scoring bias will serve as a reference for future research on mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
Evaluation Method
Figure 3: Scoring prompt template adapted from (Kim et al. 2024c), with red text indicating components that were perturbed in our experiments.
@@
Figure 2 shows our evaluation framework, inspired by (Ye et al. 2024a). On a selected dataset, we inject the scoring prompt for the LLM judges with predefined perturbation types. Then we obtain the scores from a selected judge model. We evaluate the scoring bias by calculating the correlations between judge model’s scores and golden scores, which are obtained from advanced LLMs and human annotators. Furthermore, our metrics include score distribution, which is used to measure the impact of different types of bias on the scoring tendencies of LLM judges.
@@
Table 1: Statistics of our evaluation datasets.
Problem Formulation
We assess the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and measuring the resulting impact on judgments. Following (Kim et al. 2024c), the input prompt $P$ for the scoring-based LLM judge consists of four components: a task description $T$, a target instance $C^{*}$, which can be either an instruction-response or a question-answer pair, a reference answer with a fixed score $R$, and score rubrics $S$, as shown in Figure 3. The input prompt $P$ is formulated as follows:
$P=(T,C^{*},[R],S),$ (1)
where * denotes the component will not be perturbed, and $\cdot$] means that the enclosed component is optional within the input prompt. This work focuses on biases induced by perturbations on the non-target components of the input prompt for LLM judges. Accordingly, we keep the target instance fixed. We also treat the reference answer as an optional element to explore its impact on scoring robustness, as it is not a mandatory part of the scoring process.
By perturbing the reference answer $R\rightarrow R^{\prime}$ or the score rubrics $S\rightarrow S^{\prime}$, we get a perturbed input prompt $P^{\prime}$. At the same time, the task description $T$ will also be modified to $T^{\prime}$ in conjunction with the perturbed components to ensure the prompt remains coherent. The score from the perturbed prompt will be:
$s^{\prime}=\textbf{LLM}(P^{\prime}).$ (2)
For each target instance $C$, we have a golden score $s$ from advanced LLMs or human annotators. If $s$ and $s^{\prime}$ differ, it demonstrates that the LLM judge exhibits scoring bias on the given sample.
Perturbation Types
  • Score Rubric Order: The ordering of the description for each score within the score rubrics. e.g., ascending order (score 1 to score 5), descending order (score 5 to score 1), or random order.
  • Score ID: Typically, scores are represented using Arabic numerals $\{1,2,3,4,5\}$ as IDs. In this work, we explore the use of letter grades $\{\text{E},\text{D},\text{C},\text{B},\text{A}\}$ and Roman numerals $\{\text{i},\text{ii},\text{iii},\text{iv},\text{v}\}$ as scoring IDs.
  • Reference Answer Score: Based on the finding from (Kim et al. 2024a, c) that reference answers improve scoring accuracy for LLM judges, our work addresses an unexplored question: whether attaching specific scores to these references could degrade scoring robustness.
Datasets
Following previous work (Kim et al. 2024c; Chiang, Lee, and Lukasik 2025), we select four LLM-as-a-Judge datasets for scoring bias evaluation:
  • BiGGen Bench (Kim et al. 2024b): This is a principled generation benchmark with instance-specific evaluation criteria. We use the non-multilingual portion of its test set, which contains 2,780 responses for 695 instances, all evaluated by human annotators.
  • FLASK (Ye et al. 2024b): This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
  • MT Bench (Zheng et al. 2023): This is a multi-turn chat benchmark with 80 test prompts. Kim et al. write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B (Xu et al. 2024), Vicuna-13B (Chiang et al. 2023), Llama-2-Chat-13B (Touvron et al. 2023), and GPT-3.5-Turbo-0613 (OpenAI 2023a).
  • Vicuna Bench (Chiang et al. 2023): This is a single-turn chat benchmark with 80 test prompts.Kim et al. also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
Model BiGGen Bench FLASK
@@
GPT-4.1 0.6048 0.6360 0.6401 0.6823
GPT-4o 0.6047 0.6340 0.6340 0.6557
GPT-4 0.5741 0.6045 0.6207 0.6862
Table 2: Comparison of the alignment between our LLM judgment pipeline and human golden scores, measured by Spearman’s $\rho$ and Pearson’s $r$ correlation coefficients. Scores for GPT-4.1 and GPT-4o are generated using our pipeline, whereas GPT-4 scores are sourced from the original benchmark.
The detailed statistics of our selected datasets are shown in Table 1. Due to the cost of manual annotation, the scores from advanced LLMs are usually used as golden scores (Ye et al. 2024b; Kim et al. 2024b). The instances in the above four datasets are pre-annotated with scores from GPT-4 (OpenAI 2023b). Due to the earlier release of GPT-4, we use more recent models GPT-4.1 (OpenAI 2025) and GPT-4o (OpenAI 2024) to obtain golden scores. Inspired by (Kim et al. 2024a, c), we repeat the scoring process three times using a judge prompt that includes a reference answer assigned a score of 5, and then take a majority vote to obtain the final score from the LLM evaluators. Table 2 shows that the GPT-4.1 scores exhibit the highest correlation with human scores. Throughout the remainder of our experiments, the scores generated by GPT-4.1 are treated as the golden LLM scores.
Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with a score of 5. To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4. We propose a generation-review pipeline to synthesize reference answers for each sample in FLASK, as shown in Figure 4. Given an instruction and score rubrics, the response model generates the response according to the description from score 1 to 4. Then a review model checks whether the response matches the corresponding score description. If they match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved. In this pipeline, we use GPT-4.1 and GPT-4o, alternating their roles as the response and review models.
Metrics

### v4

@@ Abstract
The “LLM-as-a-Judge” paradigm, using Large Language Models (LLMs) as automated evaluators, is pivotal to LLM development, offering scalable feedback for complex tasks. However, the reliability of these judges is compromised by various biases. Existing research has heavily concentrated on biases in comparative evaluations. In contrast, scoring-based evaluations—which assign an absolute score and are often more practical in industrial applications—remain under-investigated. To address this gap, we undertake the first dedicated examination of scoring bias in LLM judges. We shift the focus from biases tied to the evaluation targets to those originating from the scoring prompt itself. We formally define scoring bias and identify three novel, previously unstudied types: rubric order bias, score ID bias, and reference answer score bias. We propose a comprehensive framework to quantify these biases, featuring a suite of multi-faceted metrics and an automatic data synthesis pipeline to create a tailored evaluation corpus1. Our experiments empirically demonstrate that even the most advanced LLMs suffer from these substantial scoring biases. Our analysis yields actionable insights for designing more robust scoring prompts and mitigating these newly identified biases.
footnotetext: Equal contribution. Work done while at Ant Groupfootnotetext: 1Our dataset is publicly available at https://github.com/KMdsy/scoring_bias/
1 Introduction
The “LLM-as-a-Judge” [9, 15] paradigm, where Large Language Models (LLMs) serve as automated evaluators, is pivotal to LLM development. These judges provide fine-grained assessments for complex tasks like factuality evaluation [32] and instruction-following [50], and act as scalable proxies for human feedback in preference learning [2, 14]. The increasing prominence of this paradigm makes its robustness a critical concern. Ideally, LLM judges should provide objective judgments, uninfluenced by irrelevant factors. However, recent work [29] confirms that LLM judges exhibit various biases, undermining their consistency. Existing research has identified several such biases, including position bias [40], self-preference bias [42], length bias [43], and so on.
@@ 1 Introduction
While existing studies have investigated these biases, pressing issues remain. Research has heavily concentrated on comparative LLM judges, which determine the superior of two responses. However, scoring-based evaluation, providing an absolute quality score for a single response, is often more practical and widely adopted. It is more intuitive and better suited for standardized, quantitative analysis, especially when a direct comparison object is unavailable [39]. Consequently, an in-depth investigation into the biases of scoring-based judges is crucial. Research in this area remains unsystematic; while some issues like self-enhancement bias, refinement-aware bias [47], and judge model choice bias [40] have been identified, the exploration is far from comprehensive and leaves many potential biases unaddressed.
This research gap is not merely theoretical. Our work is motivated by a practical observation: when using a score-based LLM judge to evaluate a financial advisory assistant, we found that adding a single in-context example for a specific score level (e.g., an exemplar for a 4-point response) caused the absolute scores of other, unrelated samples to shift. This instability highlights a critical, unaddressed sensitivity to prompt examples and served as the primary catalyst for our investigation into other potential biases affecting score-based judges.
In this work, we consider the scoring bias as the measurable scoring shifts exhibited by an LLM-as-a-Judge when subjected to perturbations to its scoring prompt (as illustrated in Figure 1). Diverging from prior work, which has focused heavily on biases tied to the evaluation targets, our research investigates biases originating from other components of the scoring prompt. Specifically, we identify and define three novel and previously unstudied types of scoring bias: rubric order bias, score ID bias, and reference answer score bias.
To identify and quantify these three biases, we propose a comprehensive evaluation framework. This framework is designed to be extensible, enabling the identification of various scoring biases beyond those explored in this paper. We conduct experiments on four existing LLM-as-a-Judge benchmarks and utilize a suite of metrics, categorized as stability metrics, accuracy metrics, and scoring tendency, to thoroughly quantify the impact of these biases. Our key contributions can be summarized as follows:
  • We identify three novel types of scoring bias: rubric order bias, score ID bias, and reference answer score bias. To facilitate their evaluation, we introduce an automatic data synthesis pipeline that expands existing benchmarks into a tailored corpus, establishing a foundational resource for future research.
  • We propose the first evaluation framework for investigating and quantifying scoring bias, introducing metrics across three key dimensions: stability, accuracy, and scoring tendency.
  • We empirically demonstrate that even the most advanced LLMs suffer from the substantial scoring biases we identify. Furthermore, our analysis yields valuable recommendations for designing more robust scoring prompts and mitigating these biases.
2 Related Work
2.1 LLM-as-a-Judge
Previous work [9] defines LLM-as-a-Judge as the task in which an LLM generates an evaluation result in any form given input data and context. Common forms of LLM judgment include score generation [8], pairwise comparison [31], ranking [17], multiple-choice selection [16], and critique generation [10]. Improvement strategies for LLM-as-a-Judge include meticulous design of evaluation prompts [22], improvement of LLM’s abilities such as fine-tuning [41], multi-LLM collaboration pipelines such as feedback mechanisms [45], and output optimization strategies such as multi-output ensembling [34]. The performance of LLM-as-a-Judge is usually measured using metrics that represent its alignment with human annotators or advanced LLM judgments such as agreement [37], Cohen’s Kappa, and Spearman’s correlation [3]. LLM-as-a-Judge is widely applied to evaluation [19], alignment [2], retrieval [36], reasoning [6], and other specific domains [1, 33]. LLM-as-a-Judge faces challenges of bias and vulnerability, which significantly compromise the fairness and reliability of LLM judges [15].
2.2 Bias in LLM-as-a-Judge
Existing work [47] defines the prompt for LLM judges as three components: system instruction, question, and response. Based on this, they define bias in LLM-as-a-Judge as the tendency for judgments to differ when the system instruction or responses are modified in a bias-related manner. Common types of bias in LLM-as-a-Judge include position bias, length bias, self-enhancement bias, etc. These biases are often incorporated into the evaluation scope of existing LLM-as-a-Judge benchmarks [43, 51]. Currently, the types of bias remain a focal point of investigation, and numerous benchmarks have been proposed to evaluate bias in LLM-as-a-Judge. OffsetBias [29] identifies six types of bias in LLM-as-a-Judge including length bias, concreteness bias, empty reference bias, and so on. CALM [47] quantifies and analyzes 12 types of bias in LLM judges. However, the biases defined in current work are predominantly focused on comparative evaluation tasks, and there is a lack of systematic research and definitions specifically addressing biases in scoring-based evaluation tasks. Similar to research on adversarial attacks, the evaluation metrics for bias in LLM-as-a-Judge are often based on the consistency between bias-related modified judgments and the original judgments, or on alignment with golden answers.
Some studies focus on mitigating bias in LLM-as-a-Judge and aim to develop more robust LLM judges. [23] introduces Pairwise-Preference Search (PAIRS) to mitigate biases in LLM evaluators. [40] proposes three methods to calibrate positional bias, including multiple evidence calibration, balanced position calibration, and human-in-the-loop calibration. Our evaluation results on scoring bias will serve as a reference for future research on mitigating scoring biases in judge models and enhancing the robustness of scoring-based evaluations.
3 Evaluation Method
@@
Figure 2 shows our evaluation framework, inspired by [47]. On a selected dataset, we inject predefined perturbation types into the scoring prompts for the LLM judges. Then we obtain the scores from a selected judge model. We evaluate the scoring bias by calculating the consistency between scores before and after injecting perturbations, and the correlations between the judge model’s scores and the golden scores (obtained from advanced LLMs and human annotators). Furthermore, our metrics include score distribution, which is used to measure the impact of different types of bias on the scoring tendencies of LLM judges.
3.1 Problem Formulation
Scoring Prompt ### Task Description: An instruction (might include an Input inside it), a response to evaluate, a reference answer that gets a score of 5, and a score rubric representing a evaluation criteria are given. 1. Write a detailed feedback that assess the quality of the response strictly based on the given score rubric, not evaluating in general. 2. After writing a feedback, write a score that is an integer between 1 and 5. You should refer to the score rubric. 3. The output format should look as follows: “Feedback: (write a feedback for criteria) [RESULT] (an integer number between 1 and 5) 4. Please do not generate any other opening, closing, and explanations. ### Reference Answer (Score 5): {orig_reference_answer} ### Score Rubrics: [{score_criteria}] Score 1: {score1_description} Score 2: {score2_description} Score 3: {score3_description} Score 4: {score4_description} Score 5: {score5_description} ### The instruction to evaluate: {orig_instruction} ### Response to evaluate: {orig_response} ### Feedback:
Figure 3: Scoring prompt template adapted from [13], with red text indicating components that were perturbed in our experiments.
We assess the scoring bias in LLM-as-a-Judge by perturbing the scoring prompt and measuring the resulting impact on judgments. Following [13], the input prompt $P$ for the scoring-based LLM judge consists of four components: a task description $T$, an evaluation target instance $I$, which can be either an instruction-response or a question-answer pair, reference answer(s) with corresponding score(s) $A$, and score rubrics $R$ as shown in Figure 3. The input prompt $P$ is formulated as follows:
$P=(T,I^{*},[A],R),$ (1)
where * denotes the component will not be perturbed, and $\cdot$] means that the enclosed component is optional within the input prompt. This formulation is central to our study. To specifically explore biases induced by non-target components (i.e., $T$, $R$, and $[A]$), we hold the evaluation target instance $I$ constant, as denoted by $I^{*}$. We also treat the reference answer $[A]$ as an optional element to explore its impact on scoring robustness, as it is not a mandatory part of the scoring process.
By perturbing the reference answer $A\rightarrow A^{\prime}$ or the score rubrics $R\rightarrow R^{\prime}$, we get a perturbed input prompt $P^{\prime}$. At the same time, the task description $T$ will also be modified to $T^{\prime}$ in conjunction with the perturbed components to ensure the prompt remains coherent. The score from the perturbed prompt will be:
$s^{\prime}=\textbf{LLM}(P^{\prime}).$ (2)
For each target instance $I$, we have a golden score $s$ from advanced LLMs or human annotators. If $s$ and $s^{\prime}$ differ, it demonstrates that the LLM judge exhibits scoring bias on the given sample.
3.2 Perturbation Types
  • Score Rubric Order: The ordering of the description for each score within the rubric. We explore Ascending-Numeric (score 1 to score 5), Descending-Numeric (score 5 to score 1), and Random-Numeric orders (random score rubrics listed in prompt).
  • Score ID: Typically, scores are represented using Arabic numerals $\{1,2,3,4,5\}$ as IDs. In this work, we explore the use of Letter-Grades $\{\text{E},\text{D},\text{C},\text{B},\text{A}\}$ and Roman-Numerals $\{\text{i},\text{ii},\text{iii},\text{iv},\text{v}\}$ as scoring IDs.
  • Reference Answer Score: Based on the finding from [11, 13] that reference answers improve scoring accuracy for LLM judges, our work addresses an unexplored question: whether attaching a specific score $k$ to these references, denoted as Ref-$k$, could degrade scoring robustness.
3.3 Datasets
Following previous work [4, 13], we select four LLM-as-a-Judge datasets for scoring bias evaluation:
  • BiGGen Bench [12]: This is a principled generation benchmark with instance-specific evaluation criteria. We use the non-multilingual portion of its test set, which contains 2,780 responses for 695 instances, all evaluated by human annotators.
  • FLASK [48]: This is a fine-grained evaluation benchmark with 200 test prompts, 12 score rubrics, and 2,001 responses. It also includes scores labeled by human annotators.
  • MT Bench [51]: This is a multi-turn chat benchmark with 80 test prompts. [11] write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B [44], Vicuna-13B [5], Llama-2-Chat-13B [38], and GPT-3.5-Turbo-0613 [25].
  • Vicuna Bench [5]: This is a single-turn chat benchmark with 80 test prompts.[11] also write 80 hand-crafted score rubrics and generate 320 responses from WizardLM-13B, Vicuna-13B, Llama-2-Chat-13B, and GPT-3.5-Turbo-0613.
Table 1: Statistics of our evaluation datasets.
@@ 3.3 Datasets
Table 2: Comparison of the alignment between our LLM judgment pipeline and human golden scores, where $\rho$ and $r$ represent Spearman’s and Pearson’s correlation coefficients, respectively.
Model (Source of score) BiGGen Bench FLASK
@@
GPT-4.1 (our pipeline) 0.6048 0.6360 0.6401 0.6823
GPT-4o (our pipeline) 0.6047 0.6340 0.6340 0.6557
GPT-4 (original benchmark) 0.5741 0.6045 0.6207 0.6862
The detailed statistics of our selected datasets are shown in Table 1. Due to the cost of manual annotation, the scores from advanced LLMs are usually used as golden scores [12, 48]. The instances in the above four datasets are pre-annotated with scores from GPT-4 [26], which is out of date at the current time. Therefore, we use more recent models such as GPT-4.1 [28] and GPT-4o [27] to obtain golden scores. Inspired by [11, 13], we repeat the scoring process three times using a judge prompt that includes a reference answer assigned a score of 5, and then take a majority vote to obtain the final score from the LLM evaluators. Table 2 shows that the GPT-4.1 scores exhibit the highest correlation with human scores. Throughout the remainder of our experiments, the scores generated by GPT-4.1 are treated as the golden LLM scores.
Each instance in BiGGen Bench and FLASK is accompanied by a reference answer with a score of 5. To measure the impact of the reference answer score on the judgment robustness, it is necessary to additionally generate reference answers with scores ranging from 1 to 4. We propose a generation-review pipeline to synthesize reference answers for each sample in BiGGen Bench and FLASK, as shown in Figure 4. Given an instruction and score rubrics, the response model generates the response according to the description from score 1 to 4. Then a review model checks whether the response matches the corresponding score description. If they match, we accept the generated responses; otherwise, we rerun the pipeline until a match is achieved. In this pipeline, we use GPT-4.1 and GPT-4o, alternating their roles as the response and review models.
3.4 Metrics
For a dataset of $n$ samples, the human-annotated golden label set is denoted as $S=\{s_{1},\cdots,s_{n}\}$. We consider the ascending-numeric rubric without a reference answer as the baseline condition. The LLM outputs under this condition is denoted as $S^{(0)}=\{s_{1}^{(0)},\cdots,s_{n}^{(0)}\}$. Let $\mathcal{P}$ be the set of all perturbed prompt templates and the unperturbed baseline, and $\mathcal{P}^{\prime}$ be the set of all perturbed prompt templates excluding the baseline. For any prompt template $p$ in $\mathcal{P}$ or $\mathcal{P}^{\prime}$, the resulting LLM score set is $S^{(p)}=\{s_{1}^{(p)},\cdots,s_{n}^{(p)}\}$.
Stability metrics quantify the consistency between the scores generated by a perturbed prompt $p,p\in\mathcal{P}^{\prime}$ (i.e. $S^{(p)}$) and the baseline scores $S^{(0)}$. We adopt flip rate (FP) and mean absolute deviation (MAD) as metrics. FP and MAD are defined as follows,
$\text{FP}^{(p)}=\frac{1}{n}\sum_{i=1}^{n}\mathbb{I}(s_{i}^{(p)}\neq s_{i}^{(0)}),\quad\text{MAD}^{(p)}=\frac{1}{n}\sum_{i=1}^{n}|s_{i}^{(p)}-s_{i}^{(0)}|,\quad p\in\mathcal{P}^{\prime}.$ (3)
Accuracy metrics quantify the agreement between the scores generated by a perturbed prompt $p,p\in\mathcal{P}$ (i.e. $S^{(p)}$) and the golden scores $S$. Specifically, we report both Spearman’s correlation coefficient $\rho$ [35] and Pearson’s correlation coefficient $r$ [30].
Scoring tendency quantifies the model’s tendency to favor specific score categories for a given prompt template $p$ ($p\in\mathcal{P}$). For rating scale $\{r_{1},\cdots,r_{5}\}$ (e.g., $\{1,\cdots,5\}$), the tendency is defined by the following scoring distribution,
$N^{(p)}=(N_{r_{1}}^{(p)},\cdots,N_{r_{5}}^{(p)}),\quad N_{r_{k}}^{(p)}=\sum_{i=1}^{n}{\mathbb{I}(s_{i}^{(p)}=r_{k})},\quad k\in\{1,\cdots,5\}.$ (4)